OpenStudy (anonymous):

Suppose x is an angle of a right angled triangle and sin x = 12/13. tan(x)=

OpenStudy (whpalmer4):

 \[\sin x = \frac{12}{13} = \frac{\text{opposite}}{\text{hypotenuse}}\]\[\tan x = \frac{\text{opposite}}{\text{adjacent}}\]You'll need to use the Pythagorean theorem to find the side labeled ??? in order to complete the problem.